CRY+VIS Interface Meeting
Date: 2021/5/7 13:00-
Participants: Takahashi, Miyo, Tomaru, Aso, Ikeda, Ushiba, Node, Akutsu, Kimura, Hirata
- IX: IP will be released on 5/10. To adjust the length Oplev, damping of the filter chain is necessary.
- IY: Remedy of the pre-isolator is postponed.
EX: Height adjustment is on going (Status).
- EY: Installation of the tower will be restarted on 5/11. Removing of the maraging rod is necessary. Items (pure water maker, etc.) for the mirror cleaning will be ready in 2 weeks.
- Corrugate tube fixing jig: Replaced for IXV. Other 3 sets are necessary by late July.
- TM height setting: Final setting of height in EX was scheduled on 5/12. Angle adjustment in EX was scheduled on 5/13.
- Commissioning in air: Started on the pre-isolator in IX.
- Commissioning in vacuum:
- Overall commissioning: Operators are necessary.
- Acceptance check: Group work is set every Friday to make the acceptance check document.
- PF actuator: After SC.
- Jig for FC painting: Test was scheduled in the week of 5/17. Minimum target is to learn the procedure with the jig.